Bubbly:

Using the pythagorean theorem : Wilson is fishing from a small boat. His fishing hook is 7 meters below him, and a fish is swimming at the same depth as the hook, 3 meters away. How far away is Wilson from the fish? If necessary, round to the nearest tenth.

Bubbly:

6.3 or 7.6?

Bubbly:

@Gdeinward

Gdeinward:

gm bubbly

Bubbly:

good morning Gdeinward

Gdeinward:

so I figure its better to draw a digram for both you and me. I dont know if either of your answers are correct, lemme do da math

Gdeinward:



Gdeinward:



Bubbly:

so the equation for this problem should look like: a^2 + 3^2 = 7^2?

Gdeinward:

no 3^2 + 7^2 = c^2

Bubbly:

oh okay, thank you

Gdeinward:

because:

1 attachment
Bubbly:

Okay, so, the answer is 7.6 because 3^2 + 7^2= 7.61577310.. and rounded to the nearest tenth is 7.6

Gdeinward:

seems to be. You got it.

Bubbly:

thank you so much Gdeinward